OpenStudy (anonymous):

Which equations could be used to solve the problem? A 50-gallon drum is leaking water and will be empty in 12 h. At what rate is the drum being emptied? Choose exactly two answers that are correct. A. r = 12 h/50 gallons B. 50 gallons - 12 h = r C. 50 gallons - 12 h * r D. 50 gallons/12 h = r please help

OpenStudy (anonymous):

i got no idea how to solve this

OpenStudy (anonymous):

is dis rates?

OpenStudy (anonymous):

yes constant rates

OpenStudy (anonymous):

:) this is easy i did a test on this a few min ago and passed

OpenStudy (anonymous):

rly? u in k12?

OpenStudy (anonymous):

yes and answer is A.

OpenStudy (anonymous):

just divide the Hours by the gallons

OpenStudy (anonymous):

ah thanks couple more ? im rly bad at math

OpenStudy (anonymous):

ok :P but if i dont respond im afk

OpenStudy (anonymous):

wait i can pick more than one in this question do i pick another?

OpenStudy (anonymous):

or is it just A

OpenStudy (anonymous):

just a

OpenStudy (anonymous):

A and C too

OpenStudy (anonymous):

u sure?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

only A

OpenStudy (anonymous):

)))

OpenStudy (anonymous):

lol kk

OpenStudy (anonymous):

A dump truck can haul 16 loads of dirt in 4 hours. How long will the dump truck spend hauling 25 loads?

OpenStudy (anonymous):

i say 6 hours

OpenStudy (anonymous):

4 loads every 4 h

OpenStudy (anonymous):

yeah that one was ez next XD

OpenStudy (anonymous):

A pumping station pumps 340 gallons of water in 8 minutes. How many gallons does it pump in 20 minutes? A. 6800 B. 360 C. 850 D. 500

OpenStudy (anonymous):

i think its C

OpenStudy (anonymous):

brb

OpenStudy (anonymous):

k

OpenStudy (anonymous):

@iGreen can you help me?

OpenStudy (anonymous):

@ziferot do u know?

OpenStudy (anonymous):

i want to make sure that question 1 is correct with a.

OpenStudy (anonymous):

is my 3d question right ?

OpenStudy (anonymous):

340/8 * 20

OpenStudy (anonymous):

ah yes! thanks C

OpenStudy (anonymous):

Not at all

OpenStudy (amistre64):

giving out direct answers, even the wrong direct answers, is not what this site is for ....

OpenStudy (anonymous):

yeh 2. is C. thanks for helping on that one @Ziferot i was getting my food out the over xD

OpenStudy (anonymous):

The table shows a direct variation. What is the constant of variation? x y 2 4.8 4.5 10.8 6 14.4 (insert decemal here)

OpenStudy (anonymous):

@amistre64 minercity gose to k12 i also go to k12 they do things different i know how it works.

OpenStudy (amistre64):

how they work and how this website works are two different things. this is a study site ..

OpenStudy (anonymous):

true xD sorry for direct answers ill just leave now :(

OpenStudy (anonymous):

aw crud well bye =(

OpenStudy (amistre64):

question 1 is not A ....

OpenStudy (anonymous):

also i gtg :P because i got a class connect sorry @minercity

OpenStudy (anonymous):

this is my last try to get this done i just need somebody to help me solve it and ok bye =(

OpenStudy (amistre64):

we are emptying gallons per hour, so the solution will be a rate of gallons per hour.

OpenStudy (anonymous):

so um what form would that be in?

OpenStudy (amistre64):

for question 1 gallons/hour how many gallons and how many hours are stated in the question?

OpenStudy (anonymous):

50 gallons 12 h

OpenStudy (amistre64):

then 50g/12h is fine :)

OpenStudy (anonymous):

ok thanks! did my other questions get the right ancwer?

OpenStudy (amistre64):

what was the 2nd question? can you retype it?

OpenStudy (anonymous):

A dump truck can haul 16 loads of dirt in 4 hours. How long will the dump truck spend hauling 25 loads?

OpenStudy (anonymous):

i say 6 h to load 25 loads or dirt

OpenStudy (anonymous):

of*

OpenStudy (amistre64):

consider: loads per hour 16loads/4hours = 25loads/xhours 16/4 = 25/x do you know how to solve the proportion?

OpenStudy (anonymous):

4 loads pr h so 6 h to load 25

OpenStudy (anonymous):

right?

OpenStudy (amistre64):

6hrs is close; 16/4 = 4, so 4 loads an hour 6hrs*4/hr = 24 loads so we are close, but a little off.

OpenStudy (amistre64):

16/4 = 25/x 4 = 25/x 4x = 25 x = 25/4 = 6 1/4 so 6 hours and 1/4 of an hour

OpenStudy (anonymous):

what is that as a decimal? it wont fir in the box XD

OpenStudy (anonymous):

wait i know

OpenStudy (anonymous):

ok and this i have no clue what it is! The table shows a direct variation. What is the constant of variation? x y 2 4.8 4.5 10.8 6 14.4 (insert decemal here)

OpenStudy (amistre64):

a direct variation has the form: y = kx, where k is the constant of variation solving for k we get: k = y/x so pick any point, and determne the value of y/x

OpenStudy (amistre64):

4.8/2 looks simple enough

OpenStudy (anonymous):

err ok

OpenStudy (anonymous):

so 2.4?

OpenStudy (amistre64):

2.4 should be fine yes

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

6(2.4) = 14.4 so its working for us

OpenStudy (anonymous):

Which tables show a direct variation? Choose all answers that are correct. x y 1 4 3 6 5 8 A. x y 1 1.2 2 2.4 3 3.6 B. x y 5 15 10 30 15 60 C. x y 3 6 6 12 9 18 D.

OpenStudy (anonymous):

sorry its a lil messed up

OpenStudy (anonymous):

but whaterev is on top of the letter say D is the question choice

OpenStudy (anonymous):

and i can pick more than 1

OpenStudy (amistre64):

lets test them out a direct variation will have a constant multiplier, that k value x y 1 4 <<<<< notice that 1(4) = 4, so 4x is what we want 3 6 <<<<< 4(3) is not 12 so this one aint gonna be it 5 8

OpenStudy (amistre64):

x y 5 15 <<< 5k = 15, what is k gonna be? 10 30 does 10k = 30 ? 15 60 does 15k = 60 ?

OpenStudy (anonymous):

yes? and 5k

OpenStudy (amistre64):

5k = 15, 5 times 3 is 15 .. so k=3 10(3) = 30 15(3) = 45, not 60 so this cant be good can it?

OpenStudy (anonymous):

nope

OpenStudy (amistre64):

lets test the next one :)

OpenStudy (anonymous):

i think B is one of them

OpenStudy (amistre64):

we just tested B and it failed the test ...

OpenStudy (anonymous):

oh i was looking at the wrong place sorry

OpenStudy (amistre64):

if you meant C, it is good :) 3(2) = 6 6(2) =12 9(2)=18 y = 2x is a direct variation

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

hmm, you said the option LETTERs where messed up dint you

OpenStudy (anonymous):

yeah

OpenStudy (amistre64):

A x y 1 4 3 6 5 8 B x y 1 1.2 2 2.4 3 3.6 C x y 5 15 10 30 15 60 D x y 3 6 6 12 9 18

OpenStudy (anonymous):

ah there we go

OpenStudy (amistre64):

then yeah, B is good .... and D is fine as well

OpenStudy (anonymous):

kk thanks

OpenStudy (anonymous):

i was thinking when you said B was wrong i got confused XD

OpenStudy (amistre64):

yeah, these old eyes of mine had the lettering mixed up eh :)

OpenStudy (anonymous):

lol

OpenStudy (amistre64):

i have to go pick up one of my kids so ill have to let you go. becareful on this site and never accept a direct answer :) good luck

OpenStudy (anonymous):

Which equation is shown on the graph? http://static.k12.com/calms_media/media/720000_720500/720333/3/a14c2d8e51f2ee4ec464c9396b29652991304263/77003_question.jpg A y=4x B. y=x+4 C. y=1/4 D. y=x no clue

OpenStudy (anonymous):

oops didnt see your message and ok bye!

OpenStudy (amistre64):

we know y doesnt equal x since the point 10,10 is not hitting th eline we know when x=0, y doesnt equal 4 so B is out y changes so its not held constant so C is out when x=10 y = 40 so y = 4x, which is the only one left makes the most sense .. ciao :)

OpenStudy (anonymous):

thanks bye!